OpenStudy (anonymous):

solve the differential equation 2y' = y

OpenStudy (sgadi):

\[2\ln y + c\]

OpenStudy (anonymous):

say, \[y = y(t)\] \[(2/y)(dy/dt) = 1\] \[t = \int\limits (2/y)dy\]